OpenStudy (anonymous):

Evan is painting a stool that has a cylindrical seat and four rectangular prism legs. The seat has a diameter of ten inches and a height of 3 inches. Each leg is 3 inches by 3 inches by 12 inches. How much area will Evan have to paint? Use three point one four for pi. Show your work.

OpenStudy (anonymous):

@Barrelracer011

OpenStudy (barrelracer011):

Surface Area of Cylinder : = 2πrh + 2πr ² = 2 x 3.14 x 5 x 3 + 2 x 3.14 x 5 ² = 94.2 + 157 = 251.2 in. ² Area of a rectanglar prism: = 2lw + 2lh + 2wh = 2 x 3 x 3 + 2 x 3 x 12 + 2 x 3 x 12 = 18 + 72 + 72 = 162 in. ² Now multiply 162 by 4 = 648 in. ² Now add the area of the cylider and the area of the four rectangular prisms: 251.2 + 648 = 899.2 The area of the stool is 899.2 in. ²